EXPERIMENTAL CHARACTERIZATION OF IDEALIZED METAL-CENTERED REACTIONS

\underline{DARRIN BELLERT}, ADAM MANSELL, and ZACHARY THEIS, Department of Chemistry, Baylor University, Waco, TX 76798.

Instrumentation has been developed for measuring kinetic parameters for gas phase organic decompostion reactions activated by bare transition metal cations. This technique forms the reactants as a binary cluster ion in a jet-cooled expansion. A rearrangement dissociative reaction is initiated through the absorption of a visible laser photon. The temporal development of a specific fragment ion is monitored in a custom time of flight mass spectrometer. This talk focuses on the development of this instrumentation and recent results for the low-energy metal-centered ion induced decompostion reaction of organic species.